# Supplemetary Material for the paper

### v1

@@ 5 Theoretical Analysis
Let $\boldsymbol{\Lambda}=\mathbf{A}^{\top}(\mathbf{Y}-\mathbf{A}\mathbf{X})$ be the residual pre-multiplied by $\mathbf{A}^{\top}$.
@@
Lemma 5.2 (Dynamics of row norm).
Consider a system evolving under continuous gradient flow with updates as in (4), (5) and (6) with initialization $g(0)=\alpha_{g}\mathbf{1}_{L}$ and $V(0)=\alpha_{V}\mathbf{1}_{N\times L}$ where $\alpha_{g}$ and $\alpha_{V}$ are small positive scalars. Let $\mathbf{\Lambda}(t):=\mathbf{A}^{\top}(\mathbf{Y}-\mathbf{A}(\mathbf{g}(t)^{\odot 2}\mathbf{1}_{L}\odot\mathbf{V}(t)))$ and $\boldsymbol{\lambda}_{i}(t)$ denote the $i$-th row of $\mathbf{\Lambda}(t)$, and
@@
Proof sketch (for detailed proof see Lemma B.3).
@@
Lemma 5.3 (Dynamics of row norms under perfect balancedness).
@@
Proof sketch (for detailed proof see Lemma B.4).
@@
The Lemmas 5.2 and 5.3 shows that the evolution rate of each row’s norm is roughly proportional to its current norm raised to the power $4/3$. This implies that when the unbalancedness constant is small, this proportionality creates a “momentum-like” effect: smaller row norms evolve more slowly, while larger row norms change more rapidly. This dynamic, in turn, fosters an incremental learning effect: a phenomenon where different parts of the signal are learned sequentially rather than all at once. This process is realized by certain rows growing significantly in magnitude while others remain small, ultimately leading to an implicit regularization that favors low-rank solutions.
@@
Let the estimated trajectory, $\mathbf{X}(t)$ be parametrized by $({\mathbf{g}}(t),{\mathbf{V}}(t))$ and a reference trajectory $\tilde{\mathbf{X}}(t)$, be parameterized by $(\tilde{\mathbf{g}}(t),\tilde{\mathbf{V}}(t))$ both of which evolve under the continuous gradient flow, as described in Lemma C.2. Let the parameters $({\mathbf{g}}(t),{\mathbf{V}}(t))$ and $(\tilde{\mathbf{g}}(t),\tilde{\mathbf{V}}(t))$ be at a distance $D$ from the origin. At any time $t\in[0,T]$, the following inequality holds:
$\left\|\mathbf{X}(t)-\tilde{\mathbf{X}}(t)\right\|^{2}_{F}$ $\leq 8D^{4}\sum\limits_{n\in[N]}\sum\limits_{l\in[L]}\left(\left({V}_{nl}(t)-\tilde{{V}}_{nl}(t)\right)^{2}\right$.
$\quad\left.+\left({g}_{n}(t)-\tilde{g}_{n}(t)\right)^{2}\right)$ (15)
@@
This lemma provides the necessary link between the matrix trajectories and their underlying parameters. Specifically, the lemma bounds the distance between the reference trajectory, $\tilde{\mathbf{X}}(t)$, and the estimated trajectory, $\mathbf{X}(t)$ by the distance between their respective underlying parameters $({\mathbf{g}}(t),{\mathbf{V}}(t)),(\tilde{\mathbf{g}}(t),\tilde{\mathbf{V}}(t))$.
@@
Consider the system evolving under continuous gradient flow, where parameters are initialized such that perfect row-balancedness holds for all rows, i.e., $\frac{1}{2}g_{i}^{2}(0)=\sum\limits_{j\in[L]}V_{ij}^{2}(0)$ for all $i\in[N]$. Let the sensing matrix $\mathbf{A}$ have $\ell_{2}$-normalized columns and be $\mu$-coherent. Let $\mathbf{X}(t)\in\mathbb{R}^{N\times L}$ be the estimated trajectory obtained by the gradient flow with the proposed parameterization $\mathbf{X}(t)=(\mathbf{g}(t)^{\odot 2}\mathbf{1}_{L})\odot\mathbf{V}(t)$. Let $\beta$ be the smoothness constant of the loss function, $\rho$ be some constant $0<\rho\leq\alpha_{V}^{3}$, $\tilde{D}=N^{\frac{1}{3}}\sqrt{2L+1}\left(\frac{D}{2}+0.5\right)^{\frac{1}{3}}$ and $D\geq 2\rho\sqrt{K}$. If the initial value $\alpha_{V},\alpha_{g}$ (where $g(0)=\alpha_{g}\mathbf{1}_{L}$ and $V(0)=\alpha_{V}\mathbf{1}_{N\times L}$, and $\alpha_{V}=\frac{1}{\sqrt{2L}}\alpha_{g}$) is sufficiently small, specifically satisfying:
$\alpha_{V}\leq\frac{\epsilon_{app}}{2(\tilde{D}+2)^{2}}\frac{\exp(-\beta T)}{\sqrt{2LN\left({2L}+3\right)}},$ (16)
where $\epsilon_{app}\in(0,1)$, then for time $T>0$, there exists a rank-$K$ trajectory $\tilde{\mathbf{X}}(t)$ such that the estimate $\mathbf{X}(t)$ is close to $\tilde{\mathbf{X}}(t)$ until $t\geq T$ or $\|\mathbf{X}(t)\|_{F}\geq D$ i.e.,
$\|\mathbf{X}(t)-\tilde{\mathbf{X}}(t)\|_{F}<\epsilon_{app}$ (17)
for all $t\in[0,\min(T,T_{D})]$, where $T_{D}:=\inf\{t\geq 0\mid\|\mathbf{X}(t)\|_{F}\geq D\}$.
@@
$\tilde{V}_{ij}(0)$ $=\begin{cases}\rho^{1/3}\frac{V_{ij}(0)}{\|\mathbf{V}_{i:}(0)\|}\ &\forall i\in\mathcal{S}_{K}\\ 0&\forall i\notin\mathcal{S}_{K}\end{cases},\ \forall\ j\in[L],$ (19)
where, $\rho<\alpha_{V}^{3}$ is a small positive scalar. Because these non-support rows start at zero, Lemma 5.1 and Lemma 5.3, ensure they remain zero for all time $t\geq 0$. This construction guarantees the existence of a $\tilde{\mathbf{X}}(t)$, which is at most rank-$K$ matrix throughout its evolution. Note that this construction is purely for this theoretical proof; the proposed algorithm does not assume any knowledge of the support set or sparsity. To bound the distance between $\mathbf{X}(t)$ and $\tilde{\mathbf{X}}(t)$, we begin by bounding the distance between the corresponding parameters, namely, $(\mathbf{g}(t),\mathbf{V}(t))$ and $(\tilde{\mathbf{g}}(t),\tilde{\mathbf{V}}(t))$. To bound the evolution of this distance, we rely on razin2021implicit[Lemma 6]. It states that for a locally smooth objective, the distance between two gradient flow trajectories will grow at most exponentially provided the initial distance between the trajectories remains sufficiently small. Our initialization scheme (16), along with (18) and (19), ensures that at $t=0$, the distance between the parameters remains sufficiently small. Since our loss function is locally smooth, the parameter trajectories remain close.
Finally, the closeness of the parameter trajectories is translated to the distance between the trajectories $\mathbf{X}(t)$ and $\tilde{\mathbf{X}}(t)$ by applying the bound from Lemma 5.4. This leads to the theorem’s main guarantee: for a sufficiently small initialization $\alpha_{V},\alpha_{g}$, the estimated solution $\mathbf{X}(t)$ is guaranteed to remain within a user-defined distance $\epsilon_{app}$ of the reference row-sparse trajectory $\tilde{\mathbf{X}}(t)$ for a specified time duration $T$, or until the trajectory leaves a defined ball of radius $D$ around the origin. ∎
This theorem provides the theoretical guarantee for the proposed framework. It formalizes the implicit regularization effect by establishing that with a sufficiently small and balanced initialization, the algorithm’s estimated trajectory, $\mathbf{X}(t)$ is guaranteed to stay within a predefined error $(\epsilon_{app})$ of a reference rank-$K$ trajectory, $\tilde{\mathbf{X}}(t)$. By starting near the origin, the dynamics ensure that only a few rows that happen to have a slightly larger initial correlation with the true signal are selected to grow rapidly. Meanwhile balanced initialization ensures the proportional evolution of the factors required for the momentum-like dynamics to guide the algorithm’s trajectory toward the row-sparse solution.
@@
Corollary 5.1.1.
Assume the conditions of Theorem 5.1 and in addition, assume that all reference trajectories $\tilde{\mathbf{X}}(t)$ converge to a solution $\mathbf{X}^{\star}\in\mathbb{R}^{N\times L}$ and this convergence is uniform in the sense that the trajectories are confined to a bounded domain, and for any $\epsilon_{app}>0$ there exists a time $T_{c}\leq T$ after which they are all within a distance $\epsilon_{app}$ from $\mathbf{X}^{\star}$. Then for any $\epsilon_{app}>0$, if the initialization scales $\alpha_{V},\alpha_{g}$ are sufficiently small, for any time $t\in[T_{c},T]$ it holds that $\left\|\mathbf{X}(t)-\mathbf{X}^{\star}\right\|_{F}\leq 2\epsilon_{app}$.
@@
We assume that all reference trajectories $\tilde{\mathbf{X}}(t)$ converge to the oracle solution $\mathbf{X}^{\star}$ within a finite time $T>0$. This means that for any desired approximation error $\epsilon_{app}>0$ we have $\|\tilde{\mathbf{X}}(t)-\mathbf{X}^{\star}\|\leq\epsilon_{app}$.
From theorem 5.1, if the initialization values $\alpha_{V},\alpha_{g}$ are sufficiently small, the estimated trajectory is guaranteed to be within a distance of $\epsilon_{app}$ from the reference trajectory at least until time $T$. This is a conditional guarantee that holds as long as both trajectories remain within a bounded domain. We show that the estimated trajectory $\mathbf{X}(t)$} is also confined to a bounded domain. We know from our assumptions that the reference trajectory $\tilde{\mathbf{X}}(t)$ is bounded. The result follows from applying the triangle inequality. ∎
Figure 1: (a) Evolution of the norms of the components for rows in the support and non-support sets. (b) Loss versus iterations for different initialization values of $\alpha$.
(a) $\alpha_{g}=10^{-2}$
(b) $\alpha_{g}=10^{-3}$
(c) $\alpha_{g}=10^{-4}$
Figure 2: The evolution of coordinate norms during training for different initialization parameters $\alpha_{g}$.
@@ 6 Simulation
Relative Error vs M
Relative Error vs Sparsity $K$
Figure 3: Performance of IR-MMV averaged over 2020 trials.
6.1 Balancedness and Incremental Learning
@@
For this experiment, the ground truth MMV, $\mathbf{X}$, is constructed such that its active rows are set to constant values of $\mathbf{1}_{L},2\cdot\mathbf{1}_{L}$ and $3\cdot\mathbf{1}_{L}$. Recall that the parameters are initialized to ensure perfect row-wise balancedness. As illustrated in Figure 1(a), the evolution of $\frac{1}{2}g_{i}^{2}(t)$ and $\|\mathbf{V}_{i:}(t)\|_{2}^{2}$ is tracked throughout the optimization process. The empirical result support the theoretical findings of Lemma 5.1, by demonstrating that this balance is preserved over time. This dynamic leads to a momentum-like effect and an incremental learning phenomenon, where rows of the MMV are learned one at a time, a behavior consistent with prior findings by gissin2019implicit; razin2021implicit; hariz2024implicit in the context of implicit regularization via overparametrization. The magnitude of the row norms directly influences the learning speed, with smaller norms evolving more slowly and larger norms changing more rapidly. This dynamic implicitly promotes structured sparsity. The time-invariant balancedness property (Lemma 5.1) guarantees that if an element $g_{i}(t)$ approaches zero, the corresponding row $\mathbf{V}_{i:}$ will also converge to zero, rendering the entire $i$-th row of $\mathbf{X}$ sparse.
6.2 Effect of Initialization
To analyze the effect of initialization, the experiment is repeated for a set of different initializations, specifically $\alpha_{g}\in\{10^{-2},10^{-3},10^{-4}\}$ and the corresponding $\alpha_{V}$. All other parameters, are kept constant across these runs. For this experiment, the ground truth MMV, $\mathbf{X}$, is generated as in the previous section. The simulation results highlight the importance of initialization values. As seen in Figure 1(b), the convergence behavior of the loss function varies significantly with the choice of $\alpha_{g}$. It demonstrates that smaller initialization values lead to convergence to a solution with a lower final loss. This empirical finding is consistent with Theorem 5.1, which formally establishes that a smaller initialization value ($\alpha_{V}$) is necessary to guarantee a smaller approximation error ($\epsilon_{app}$). Furthermore, Figure 2 illustrates how growth of coordinate norms is influenced by the initialization. A smaller value results in a slower, more gradual convergence, while larger initial values may lead to faster but less stable convergence.
6.3 Performance with Varying Noise
The empirical performance of IR-MMV is compared against different MMV reconstruction algorithms (M-OMP, M-SP and M-FOCUSS) for various values of $M$ and $K$. For these experiments, the measurement matrix $\mathbf{Y}$ is generated by adding noise to the signal $(\mathbf{Y}=\mathbf{A}\mathbf{X}+\mathbf{W})$, such that the Signal-to-Noise Ratio is 4040dB. Performance analysis is evaluated using relative estimation error $e_{r}=\frac{\|\mathbf{X}-\hat{\mathbf{X}}\|_{F}^{2}}{\|\mathbf{X}\|_{F}^{2}}$. The results, averaged over $20$ trials, are shown in Figure 3; the shaded regions represent the standard deviation across the trials. The proposed method demonstrates performance comparable to the other approaches. Critically, it achieves this without any parameter tuning or prior knowledge of data-specific parameters. In contrast, the benchmark algorithms require such information to perform optimally; both M-OMP and M-SP requires prior knowledge of true sparsity level $K$. M-FOCUSS requires proper tuning of a parameter $\lambda$ and/or knowledge of noise variance. We defer the results on MNIST data to Appendix D.

### v2

@@ 5 Theoretical Analysis
We begin by identifying a key conservation law of the optimization dynamics for the proposed approach. The following lemma shows that the proposed overparameterization preserves a balancedness property, both globally and at the row level, throughout the optimization process.
@@
In the general case, $\epsilon$ and $\epsilon_{r}$ are independent quantities. However, under the specific initialization of Algorithm 1 (where $g(0)$ and $V(0)$ are initialized with constant values $\alpha_{g}$ and $\alpha_{V}$), it strictly holds that $\epsilon_{r}\leq\epsilon$ for all $t\geq 0$. (See Appendix B for the proof.)
We now characterize how the norms of individual rows of the reconstructed signal evolve under gradient flow. The next lemma provides upper and lower bounds on the rate of change of each row norm.
Theorem 5.3 (Dynamics of row norm).
Consider a system evolving under continuous gradient flow with updates equations (4), (5) and (6) with initialization $g(0)=\alpha_{g}\mathbf{1}_{L}$ and $V(0)=\alpha_{V}\mathbf{1}_{N\times L}$ where $\alpha_{g}$ and $\alpha_{V}$ are small positive scalars. Let $\mathbf{\Lambda}(t):=\mathbf{A}^{\top}(\mathbf{Y}-\mathbf{A}(\mathbf{g}(t)^{\odot 2}\mathbf{1}_{L}\odot\mathbf{V}(t)))$ and $\boldsymbol{\lambda}_{i}(t)$ denote the $i$-th row of $\mathbf{\Lambda}(t)$, and
@@
Proof sketch (for detailed proof see Theorem B.4).
@@
To gain further insight into the row-wise dynamics, we next consider the special case of perfect row-balancedness. Under this condition, the evolution of each row norm admits a simplified and more explicit characterization.
Corollary 5.4 (Dynamics of row norms under perfect balancedness).
@@
Proof sketch (for detailed proof see Corollary B.5).
@@
Theorem 5.3 and Corollary 5.4 show that the evolution rate of each row’s norm is roughly proportional to its current norm raised to the power $4/3$. This implies that when the unbalancedness constant is small, this proportionality creates a “momentum-like” effect: smaller row norms evolve more slowly, while larger row norms change more rapidly. This dynamic, in turn, fosters an incremental learning effect: a phenomenon where different parts of the signal are learned sequentially rather than all at once. This process is realized by certain rows growing significantly in magnitude while others remain small, ultimately leading to an implicit regularization that favors low-rank solutions. The following lemma connects the behavior of the optimization in parameter space to the evolution of the reconstructed signal. It shows that the distance between two signal trajectories can be upper-bounded by the distance between their corresponding parameters.
@@
Let the estimated trajectory, $\mathbf{X}(t)$ be parametrized by $({\mathbf{g}}(t),{\mathbf{V}}(t))$ and a reference trajectory $\tilde{\mathbf{X}}(t)$, be parameterized by $(\tilde{\mathbf{g}}(t),\tilde{\mathbf{V}}(t))$ both of which evolve under the continuous gradient flow, as described in Lemma 5.1. Let $T>0$ and $D>0$ be scalars. Specifically, assume that for all $t\in[0,T]$ and all $n\in[N]$, the condition ${\max}\left\{|g_{n}(t)|,\|V_{n:}(t)\|_{2},|\tilde{g}_{n}(t)|,\|\tilde{V}_{n:}(t),\|_{2}\right\}\leq D$ holds. Then at any time $t\in[0,T]$, the following inequality holds:
$\left\|\mathbf{X}(t)-\tilde{\mathbf{X}}(t)\right\|^{2}_{F}$ $\leq 8D^{4}\Big\|(\mathbf{g}(t),\mathbf{V}(t))\left.-(\tilde{\mathbf{g}}(t),\tilde{\mathbf{V}}(t))\right\|^{2},$ (15)
where, $\Big\|(\mathbf{g}(t),\mathbf{V}(t))\left.-(\tilde{\mathbf{g}}(t),\tilde{\mathbf{V}}(t))\right\|^{2}=\sum\limits_{n\in[N]}\sum\limits_{l\in[L]}\left(\left({V}_{nl}(t)-\tilde{{V}}_{nl}(t)\right)^{2}\right.\left.+\left({g}_{n}(t)-\tilde{g}_{n}(t)\right)^{2}\right)$.
@@
This provides the necessary link between the matrix trajectories and their underlying parameters, thus serving as the final component required for the convergence analysis. Specifically, the lemma bounds the distance between the reference trajectory, $\tilde{\mathbf{X}}(t)$, and the estimated trajectory, $\mathbf{X}(t)$ by the distance between their respective underlying parameters $({\mathbf{g}}(t),{\mathbf{V}}(t)),(\tilde{\mathbf{g}}(t),\tilde{\mathbf{V}}(t))$. Building on this analysis, the following theorem establishes our main convergence guarantee.
@@
Consider the system evolving under continuous gradient flow, where parameters are initialized such that perfect row-balancedness holds for all rows, i.e., $\frac{1}{2}g_{i}^{2}(0)=\sum\limits_{j\in[L]}V_{ij}^{2}(0)$ for all $i\in[N]$. Let the sensing matrix $\mathbf{A}$ have $\ell_{2}$-normalized columns and be $\mu$-coherent. Let $\mathbf{X}(t)\in\mathbb{R}^{N\times L}$ be the estimated trajectory obtained by the gradient flow with the proposed parameterization $\mathbf{X}(t)=(\mathbf{g}(t)^{\odot 2}\mathbf{1}_{L})\odot\mathbf{V}(t)$. If the initial value $\alpha_{V},\alpha_{g}$ (where $g(0)=\alpha_{g}\mathbf{1}_{L}$ and $V(0)=\alpha_{V}\mathbf{1}_{N\times L}$, and $\alpha_{V}=\frac{1}{\sqrt{2L}}\alpha_{g}$) is sufficiently small, specifically satisfying:
${\alpha_{V}}$ $\leq\frac{\epsilon_{\text{app}}}{{4D^{2}}{}\sqrt{2NL(2L+1)}},$ (16)
where $\epsilon_{\text{app}}\in(0,1)$, then for time $T>0$, there exists a rank-$K$ trajectory $\tilde{\mathbf{X}}(t)$ such that the estimate $\mathbf{X}(t)$ is close to $\tilde{\mathbf{X}}(t)$ until $t\geq T$, i.e.,
$\|\mathbf{X}(t)-\tilde{\mathbf{X}}(t)\|_{F}<\epsilon_{\text{app}}$ (17)
for all $t\in[0,T]$.
@@
$\tilde{V}_{ij}(0)$ $=\begin{cases}\frac{1}{\sqrt{L}}\rho^{1/3}\ &\forall i\in\mathcal{S}_{K}\\ 0&\forall i\notin\mathcal{S}_{K}\end{cases},\ \forall\ j\in[L],$ (19)
where, $\rho<L^{3/2}\alpha_{V}^{3}$ is a small positive scalar. Because these non-support rows for reference trajectory start at zero, Lemma 5.1 and Corollary 5.4, ensure they remain zero for all time $t\geq 0$. This construction guarantees the existence of a $\tilde{\mathbf{X}}(t)$, which is at most rank-$K$ matrix throughout its evolution. Note that this construction is purely for this theoretical proof; the proposed algorithm does not assume any knowledge of the support set or sparsity. To bound the distance between $\mathbf{X}(t)$ and $\tilde{\mathbf{X}}(t)$, we analyze the divergence of their corresponding parameters using a Lyapunov function, $\mathcal{E}(t)=\Big\|\big(\mathbf{g}(t),\mathbf{V}(t)\big)-\big(\tilde{\mathbf{g}}(t),\tilde{\mathbf{V}}(t)\big)\Big\|^{2}$. This allows practical initialization scale as is apparent in Equation (16). This stands in contrast to standard analysis techniques that rely on Grönwall’s inequality (Razin et al., 2021; Gidel et al., 2019), which will lead to an initialization upper bound which will be exponentially smaller in $T$ (i.e., $\exp(-\beta T),$ where $\beta$ is the smoothness constant).
The Lyapunov function is decomposed into three components: the energy of the estimated parameters, the energy of reference parameters and the corresponding cross term. The time derivative of each term is then upper bounded with a time-integrable growth. Integrating these bounds yields a rational polynomial-in-time bound on the Lyapunov function, ensuring that the parameter distance remains below a prescribed threshold for all $t\in[0,T]$ provided the initialization satisfies (16).
Finally, the closeness of the parameter trajectories is translated to the distance between the trajectories $\mathbf{X}(t)$ and $\tilde{\mathbf{X}}(t)$ by applying the bound from Lemma 5.5. This leads to the theorem’s main guarantee: for a sufficiently small initialization $\alpha_{V},\alpha_{g}$, the estimated solution $\mathbf{X}(t)$ is guaranteed to remain within a user-defined distance $\epsilon_{\text{app}}$ of the reference row-sparse trajectory $\tilde{\mathbf{X}}(t)$ for a specified time duration.∎
This theorem provides the theoretical guarantee for the proposed framework. It formalizes the implicit regularization effect by establishing that with a sufficiently small and balanced initialization, the algorithm’s estimated trajectory, $\mathbf{X}(t)$ is guaranteed to stay within a predefined error $(\epsilon_{\text{app}})$ of a reference rank-$K$ trajectory, $\tilde{\mathbf{X}}(t)$. By starting near the origin, the dynamics ensure that only a few rows that happen to have a slightly larger initial correlation with the true signal are selected to grow rapidly. Meanwhile balanced initialization ensures the proportional evolution of the factors required for the momentum-like dynamics to guide the algorithm’s trajectory toward the row-sparse solution.
@@
Corollary 5.7.
Assume the conditions of Theorem 5.6 and in addition, assume that all reference trajectories $\tilde{\mathbf{X}}(t)$ converge to a solution $\mathbf{X}^{\star}\in\mathbb{R}^{N\times L}$ and this convergence is uniform in the sense that the trajectories are confined to a bounded domain, and for any $\epsilon_{\text{app}}>0$ there exists a time $T_{c}\leq T$ after which they are all within a distance $\epsilon_{\text{app}}$ from $\mathbf{X}^{\star}$. Then for any $\epsilon_{\text{app}}>0$, if the initialization scales $\alpha_{V},\alpha_{g}$ are sufficiently small, for any time $t\in[T_{c},T]$ it holds that $\left\|\mathbf{X}(t)-\mathbf{X}^{\star}\right\|_{F}\leq 2\epsilon_{\text{app}}$.
@@
We assume that all reference trajectories $\tilde{\mathbf{X}}(t)$ converge to the oracle solution $\mathbf{X}^{\star}$ within a finite time $T>0$. This means that for any desired approximation error $\epsilon_{\text{app}}>0$ we have $\|\tilde{\mathbf{X}}(t)-\mathbf{X}^{\star}\|\leq\epsilon_{\text{app}}$.
From Theorem 5.6, if the initialization values $\alpha_{V},\alpha_{g}$ are sufficiently small, the estimated trajectory is guaranteed to be within a distance of $\epsilon_{\text{app}}$ from the reference trajectory at least until time $T$. This is a conditional guarantee that holds as long as both trajectories remain within a bounded domain. We show that the estimated trajectory $\mathbf{X}(t)$} is also confined to a bounded domain. We know from our assumptions that the reference trajectory $\tilde{\mathbf{X}}(t)$ is bounded. The result follows from applying the triangle inequality. ∎
@@ 6 Simulation
In this section, we present simulations to reinforce our theoretical results. We also compare the performance of our proposed IR-MMV approach against five established MMV recovery algorithms: M-OMP, M-SP, AMP-MMV, M-BSBL and M-FOCUSS. Unless otherwise specified, the default values for simulation parameters are: $M=500$, $N=10000$, $L=20$, row sparsity level $K=3$ $(K<N)$ typically used in IR/MMV literature. The sensing matrix $\mathbf{A}$ is generated by drawing its entries from an i.i.d. sub-Gaussian distribution (to satisfy $\mu$-coherence property) with $\ell_{2}$-normalized columns. In line with our theoretical analysis, the algorithm’s hyperparameters are configured as follows: $\mathbf{g}$ is initialized with $\alpha_{V}=5\times 10^{-4}$ to satisfy the small initialization requirement of Theorem 5.6, while the component matrix $\mathbf{V}$ is initialized to $\alpha_{V}=\frac{1}{\sqrt{2L}}\alpha_{g}$ to ensure the perfect row-wise balancedness. Finally, the maximum number of iterations is set to $T=5\times 10^{6}$ to provide the algorithm with sufficient time to converge from its deliberately small initial state. For our experiments, the parameters of M-FOCUSS are set to $p=0.8$ and $\lambda=0.01$ as recommended by Cotter et al. (2005). Additionally, the exact sparsity level $K$ was provided to M-OMP, M-SP and AMP-MMV, with the latter also receiving the true noise variance.
6.1 Balancedness and Incremental Learning
@@
This dynamic leads to a momentum-like effect and an incremental learning phenomenon, where rows of the MMV are learned one at a time, a behavior consistent with prior findings by Gissin et al. (2019); Razin et al. (2021); Hariz et al. (2024) in the context of implicit regularization via overparametrization. The magnitude of the row norms directly influences the learning speed, with smaller norms evolving more slowly and larger norms changing more rapidly. This dynamic implicitly promotes structured sparsity. The time-invariant balancedness property (Lemma 5.1) guarantees that if an element $g_{i}(t)$ approaches zero, the corresponding row $\mathbf{V}_{i:}$ will also converge to zero, rendering the entire $i$-th row of $\mathbf{X}$ sparse.
Figure 1: Evolution of the norms of the components for rows in the support and non-support sets
Figure 2: Performance comparison of the proposed IR-MMV approach against baseline algorithms with both optimal and mismatched parameters. (a) Support Recovery (F1-Score) vs. SNR (b) Relative Error (RMSE) vs. SNR, (c) Support Recovery (F1-Score) vs. Sparsity $(K)$ and (d) Relative Error (RMSE) vs. Sparsity $(K)$.
6.2 Performance Comparison with Other Algorithms
[…]
As seen in Figure 2, our approach consistently achieves performance comparable to the optimally tuned baselines. Our approach achieves this performance without any parameter tuning or prior knowledge of data-specific parameters. While M-BSBL does not require sparsity priors, they exhibit poor performance in the low SNR regime (0(0dB - 1010dB)). This is likely because, in these high-noise settings, reliably learning the hyperparameters from the data becomes infeasible. We also observe that the RMSE of M-FOCUSS remains consistently higher than that of the other approaches, even though its F11 score is $1$.
The performance of M-OMP, M-SP, M-FOCUSS, and AMP-MMV drops when we introduce parameter mismatches (e.g., $K\pm 1$, $\lambda\pm 10\%$ and $\pm 10\%$ deviation in noise variance). As observed in Figures 2(b), 2(d), the greedy methods (M-OMP, M-SP) perform very poorly when the sparsity level is underestimated (K1K-1), failing to converge even at high SNR. When the sparsity is overestimated ($K+1$), the performance deteriorates, though the degradation is less severe compared to the underestimation case. In contrast, overestimation yields better error rates but prevents these algorithms from achieving perfect support recovery, with F11-scores saturating below $1$ (Figures 2(a), 2(c)). We defer the results on MNIST data to Appendix D.
Summarizing, methods such as M-OMP, M-SP, M-FOCUSS, and AMP-MMV degrade in the presence of parameter misspecification while our method does not. Furthermore, we also outperform Bayesian approaches like M-BSBL at low SNR. Hence, our approach has significant utility in practical scenarios where typically there is parameter misspecification and/or low SNR.